OpenStudy (anonymous):

what is the area of a triangle whos base is 6 2/9 and height is 2 6/11?

OpenStudy (anonymous):

(6+2/9)(2+6/11)/2= 56/9*28/11*1/2= 28*28/99= 784/99= 7.(91)

jhonyy9 (jhonyy9):

A=(b*h)/2

OpenStudy (anonymous):

i know that

jhonyy9 (jhonyy9):

so than make the calcule

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

i did it